OpenStudy (anonymous):

3m/4+9n=-2(4n/5-6m/5) solve for m

jhonyy9 (jhonyy9):

(3m+36n)/4 =-2(4n-6m)/5 5(3m+36n) = -8(4n-6m) 15m+180n = -32n +48m 212n=33m m= 212n/33